# A Survey on Safe Multi-Modal Learning System

### 08 Feb 2024 v1

@@
Abstract
With the wide deployment of multimodal learning systems (MMLS) in real-world scenarios, safety concerns have become increasingly prominent. The absence of systematic research into their safety is a significant barrier to progress in this field. To bridge the gap, we present the first taxonomy for MMLS safety, identifying four essential pillars of these concerns. Leveraging this taxonomy, we conduct in-depth reviews for each pillar, highlighting key limitations based on the current state of development. Finally, we pinpoint unique challenges in MMLS safety and provide potential directions for future research.
1 Introduction
@@
Multimodal learning has rapidly evolved in the past few years. Compared to unimodal learning, it could more closely mimic human information processing, where multiple modalities, such as text, image, and audio, are incorporated to understand and interact with the world Baltrušaitis et al. (2018).
@@
Despite significant advancements in related research areas, deploying multimodal learning systems (MMLS) in real-world situations, especially in high-stakes environments like healthcare and autonomous driving, naturally brings about safety challenges. These challenges are notably more intricate and acute than in unimodal environments. Firstly, when implemented in real-world situations, MMLS are likely to face various types of uncommon events and risks. These include natural distribution shifts and deliberately crafted attacks intended to disrupt the systems. The complexity introduced by the multimodal setting may exacerbate these concerns, as it not only involves mismatches within each modality but also introduces domain shifts across modalities McKinzie et al. (2023). Moreover, the rich information spanned by multiple modalities opens doors for cross-modal attacks by malicious entities, further increasing MMLS vulnerability Zhang et al. (2022). The growing risk of privacy is another significant concern. The integration of multiple modalities enhances model efficacy but simultaneously provides attackers with additional information like modality alignment details Rao et al. (2023), exacerbating breaches of privacy. Furthermore, the increasing scales of MMLS further aggravate the challenge of pre-trained models inadvertently memorizing sensitive information from their training data Naseh et al. (2023). MMLS are also at high risk of losing control and generating harmful responses such as discriminatory information or misinformation Schlarmann and Hein (2023); Li et al. (2024). Due to their increasing complexity, especially in large-scale MMLS, monitoring these systems become more challenging. This increased difficulty in oversight makes them more vulnerable to manipulation, resulting in a greater likelihood of performing harmful actions compared to their unimodal counterparts.
Figure 1: Taxonomy for Multimodal Learning Safety.
@@
To advance the development of MMLS, it is essential to thoroughly consider these unique safety challenges and develop appropriate strategies. This survey serves to review recent advances in safe MMLS from four primary pillars: robustness, privacy, monitoring, and controllability. For robustness, we explore two main aspects: the resilience of MMLS against distribution shifts, and their robustness in the face of adversarial perturbations and attacks. Regarding privacy, we first highlight some unique privacy challenges inherent in MMLS, followed by a discussion on various paradigms for preserving privacy in these contexts. For monitoring, our focus is twofold. First, we examine anomaly detection, which is crucial for promptly identifying potential hazards. Second, we investigate the reliability of model outputs to enhance the trustworthiness of MMLS. Controllability is divided into two subcategories. The first, interpretability, aims to study and increase the transparency and understandability of MMLS. The second, fairness, addresses inherent bias in these systems to ensure more equitable and fair responses. We also summarize pertinent benchmarks related to these pillars and conclude the paper with open challenges.
@@
In summary, our main contributions are Taxonomy: We present the first taxonomy for MMLS safety; Review: Following this new taxonomy, we conduct a comprehensive review of existing works, highlighting their limitations; and Open Challenges: We identify some unique challenges related to the safety of MMLS and suggest potential prospects.
2 Robustness
@@
2.1 Robustness against Distributional Shift
@@
2.1.1 Understanding distributional robustness in MMLS
@@
Most prior research on robust MMLS focuses on vision-language models (VLMs), which are susceptible to distribution shifts in both image and text modalities. Yet they generally show greater robustness than unimodal models Fang et al. (2022). Understanding the reasons behind this robustness gain is vital for advancing MMLS safety. For instance, Fang et al. (2022) identifies diverse image training data as the key to CLIP’s robustness, excluding language supervision and contrastive loss function. However, contrasting views exist Xue et al. (2023), suggesting these latter factors also make contributions. Consensus on these findings is yet to be established. Modality mismatch represents a unique distribution shift in MMLS, which occurs when the data modalities during testing differ from those in training. Earlier studies primarily explored the impact of missing modalities on MMLS’s robustness Ma et al. (2022). More recent research starts considering more generalizable settings McKinzie et al. (2023), investigating the effects of both missing and newly added modalities during test time on MMLS robustness.
2.1.2 Improving MMLS distributional robustness
@@
Current efforts to improve the distributional robustness of MMLS often involve employing data augmentation techniques and implementing robust training strategies.
@@
Data Augmentation. Data augmentation techniques, proven effective in unimodal systems for enhancing model robustness, are naturally extended to MMLS. However, applying unimodal strategies independently to each modality risks losing semantic coherence Liu et al. (2022). Consequently, multimodal techniques preserving modality consistency while improving robustness have emerged. For example, RobustMixGen for VLMs uses a pre-synthesis data separation module to separate data into objects and backgrounds. It maintains cross-modal semantics and boosts model robustness by reducing spurious correlations Kim et al. (2023).
@@
Robust Training. Research in robust training strategies is diverse and spans from improving pre-training strategies to adaptation and fine-tuning methods. For instance, a multi-task optimization algorithm has been developed to automatically identify the optimal data fusion strategies, enhancing models’ robustness against cross-modal distribution shifts Ma et al. (2022). Andonian et al. (2022) presents a novel contrastive learning method for pretraining, combining knowledge distillation and soft alignments to improve robustness. Other studies include devising robust adaptation methods. For instance, Wortsman et al. (2022) shows that fusing weights from zero-shot and fine-tuned CLIPs enhances distributional robustness, with further improvements observed through weight averaging from CLIPs fine-tuned under different hyperparameters.
2.2 Adversarial Robustness
@@
MMLS robustness to adversarial attacks faces unique challenges. Malicious users might use the information from multiple modalities to breach security. Cross-modal adversarial attacks, such as subtle alterations in an image affecting text predictions, further complicate the issue.
2.2.1 Multimodal adversarial attack.
@@
Multimodal adversarial attack aims to trick a target MMLS by subtly altering its multimodal inputs. The objective function for multimodal adversarial attacks can be formulated as:
$\small\mathop{\arg\max}_{\{x_{i}^{adv}\}_{i=1}^{n}}\mathcal{L}(\{x_{i}^{adv}\}_{i=1}^{n},y;\theta)\quad s.t.\,||x_{i}^{adv}-x_{i}||_{p}\leq\epsilon_{i}$ (1)
@@
where $\{x_{i}\}_{i=1}^{n}$ denotes the input composed of $n$ modalities, $\{x_{i}^{adv}\}_{i=1}^{n}$ are the generated adversarial samples, and $\epsilon_{i}$ is the perturbation budget for the $i$-th modality. The attacker optimizes $\mathcal{L}$ to induce model mispredictions within permissible budgets. An effective multimodal adversarial attack needs to properly address the modality gap between multiple input data modalities. Zhang et al. (2022) studies multimodal attacks on VLMs, discovering that launching unimodal attacks independently often fails. They then propose Co-Attack, which coordinates image and text attacks through a sequential strategy Another challenge lies in executing effective, downstream-agnostic attacks, especially within the prevalent ’pretrain-then-finetune’ framework for MMLS, as designing specific attacks for each task a pre-trained encoder might tackle is impractical. Pioneering work by Zhou et al. (2023b) proposes a method for generating broadly applicable adversarial noise, ensuring adversarial samples’ significant deviation from decision boundary in the feature space via a topology-deviation-based generative adversarial network.
2.2.2 Enhance adversarial robustness of multimodal models.
@@
Current works on improving the adversarial robustness of MMLS predominantly center around robust fusion strategy and adversarial training.
@@
Robust Fusion and Alignment. The unique modality alignment process of MMLS, compared to unimodal models, has catalyzed extensive research into enhancing their adversarial robustness. For instance, Yang et al. (2021) improves MMLS single-source adversarial robustness by designing a fusion strategy that detects inconsistencies across modalities and executes feature fusion to counteract the perturbed modality. A constrastive learning framework has been proposed that captures both intra- and inter-modality dynamics, filtering out noise present in various modalities and thereby yielding more robust representations Lin and Hu (2022).
@@
Adversarial Training. Adversarial training enhances robustness effectively, yet its application in MMLS presents challenges, including higher computational demands and scalability issues, compared to unimodal settings. To approach these, Gan et al. (2020) adopts the free adversarial training strategy that obtains parameters’ gradients with almost no extra computational expense. Another challenge lies in improving the generalization of adversarial training procedures within the emerging pretrain-then-finetune paradigm for MMLS. An adversarial pre-training strategy for VLMs is introduced in Gan et al. (2020), focusing on adversarial training in the embedding space of each modality rather than directly perturbing image pixels or text tokens, effectively enhancing model performance in a task-agnostic manner.
2.3 Benchmark
@@
The assessment of model distributional robustness commonly employs metrics such as Effective Robustness, along with Relative Robustness and the MultiModal Impact Score to measure average performance drops during distributional shifts Fang et al. (2022); Qiu et al. (2022). Systematic benchmarks are constantly being proposed as well. For instance, Li et al. (2020) integrates 9 VQA datasets to study four generic robustness in VLMs. Recently, benchmarks for large MMLS have emerged, such as the one introduced in Han et al. (2023). This benchmark assesses GPT-4V’s zero-shot generalization on 13 varied datasets, offering insights into its adaptability to distribution shifts.
2.4 Discussion

### 22 Mar 2024 v2

@@
Abstract.
In the rapidly evolving landscape of artificial intelligence, multimodal learning systems (MMLS) have gained traction for their ability to process and integrate information from diverse modality inputs. Their expanding use in vital sectors such as healthcare has made safety assurance a critical concern. However, the absence of systematic research into their safety is a significant barrier to progress in this field. To bridge the gap, we present the first taxonomy that systematically categorizes and assesses MMLS safety. This taxonomy is structured around four fundamental pillars that are critical to ensuring the safety of MMLS: robustness, alignment, monitoring, and controllability. Leveraging this taxonomy, we review existing methodologies, benchmarks, and the current state of research, while also pinpointing the principal limitations and gaps in knowledge. Finally, we discuss unique challenges in MMLS safety. In illuminating these challenges, we aim to pave the way for future research, proposing potential directions that could lead to significant advancements in the safety protocols of MMLS.
Multi-modal System, AI Safety, Responsible AI
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Multimodal learning has rapidly evolved across various domains in the past few years. Compared to unimodal learning, it could more closely mimic human information processing, where multiple modalities, such as text, image, and audio, are incorporated to understand and interact with the world (Baltrušaitis et al., 2018).
@@
Despite significant advancements in related research areas, deploying multimodal learning systems (MMLS) in real-world situations, especially in high-stakes environments like healthcare and autonomous driving, naturally brings safety challenges. These challenges are notably more intricate and acute than in unimodal environments (Han et al., 2022). Firstly, when implemented in real-world situations, MMLS are likely to face various types of uncommon events and risks. These include natural distribution shifts and deliberately crafted attacks intended to disrupt the systems. The complexity introduced by the multimodal setting may exacerbate these concerns, as it not only involves mismatches within each modality but also introduces domain shifts across modalities (McKinzie et al., 2023). Moreover, the rich information spanned by multiple modalities opens doors for cross-modal attacks by malicious entities, further increasing MMLS vulnerability (Zhang et al., 2022). The growing risk of privacy is another significant concern. The integration of multiple modalities enhances model efficacy but simultaneously provides attackers with additional information like modality alignment details (Rao et al., 2023), exacerbating breaches of privacy. Furthermore, the increasing scales of MMLS further aggravate the challenge of pre-trained models inadvertently memorizing sensitive information from their training data (Naseh et al., 2023). MMLS are also at high risk of losing control and generating harmful responses such as discriminatory information or misinformation (Schlarmann and Hein, 2023; Li et al., 2024). Due to their increasing complexity, especially in large-scale MMLS, monitoring these systems becomes more challenging. This increased difficulty in oversight makes them more vulnerable to manipulation, resulting in a greater likelihood of performing harmful actions compared to their unimodal counterparts.
Figure 1. Taxonomy for Safety of Multimodal Learning Systems.
@@
To advance the development of MMLS, it is essential to thoroughly consider these unique safety challenges and develop appropriate strategies. This survey serves to review recent advances in safe MMLS from four primary pillars: robustness, alignment, monitoring, and controllability. For robustness, we explore two main aspects: the resilience of MMLS against distribution shifts, and their robustness in the face of adversarial perturbations and attacks. Regarding alignment, we first examine the issue of the misalignment phenomenon in MMLS, noting their susceptibility to being jailbroken and producing harmful content. We then proceed to introduce the current techniques developed for aligning MMLS. For monitoring, our focus is twofold. First, we examine anomaly detection, which is crucial for promptly identifying potential hazards. Second, we investigate the reliability of model outputs to enhance the trustworthiness of MMLS. Controllability is divided into three subcategories. The first, interpretability, aims to study and increase the transparency and understandability of MMLS. The second, privacy, discusses unique privacy challenges within MMLS, and various paradigms for preserving privacy in these contexts. Lastly, fairness addresses inherent bias in these systems to ensure more equitable and fair responses. We also discuss the limitations and summarize pertinent benchmarks related to these pillars and conclude the paper with open challenges.
@@
In summary, our main contributions are Taxonomy: We present the first taxonomy (Figure 1) for MMLS safety; Review: Following this new taxonomy, we conduct a comprehensive review of existing works, highlighting their limitations; and Open Challenges: We identify some unique challenges related to the safety of MMLS and suggest potential prospects.
2. Robustness
@@
2.1. Robustness against Distributional Shift
@@
2.1.1. Understanding distributional robustness in MMLS
@@
Most prior research on robust MMLS focuses on vision-language models (VLMs), which are susceptible to distribution shifts in both image and text modalities. Yet they generally show greater robustness than unimodal models (Fang et al., 2022). Understanding the reasons behind this robustness gain is vital for advancing MMLS safety. For instance, (Fang et al., 2022) identifies diverse image training data as the key to CLIP’s robustness, excluding language supervision and contrastive loss function. However, contrasting views exist (Xue et al., 2023), suggesting these latter factors also make contributions. Consensus on these findings is yet to be established. Modality mismatch represents a unique distribution shift in MMLS, which occurs when the data modalities during testing differ from those in training. Earlier studies primarily explored the impact of missing modalities on MMLS’s robustness (Ma et al., 2022). More recent research starts considering more generalizable settings (McKinzie et al., 2023), investigating the effects of both missing and newly added modalities during test time on MMLS robustness.
2.1.2. Improving MMLS distributional robustness
@@
Current efforts to improve the distributional robustness of MMLS often involve data augmentation techniques and robust training strategies.
@@
Data Augmentation. Data augmentation techniques, proven effective in unimodal systems for enhancing model robustness, are naturally extended to MMLS. However, applying unimodal strategies independently to each modality risks losing semantic coherence (Liu et al., 2022). Consequently, multimodal techniques preserving modality consistency while improving robustness have emerged. For example, RobustMixGen for VLMs uses a pre-synthesis data separation module to separate data into objects and backgrounds. It maintains cross-modal semantics and boosts model robustness by reducing spurious correlations (Kim et al., 2023).
@@
Robust Training. Research in robust training strategies is diverse and spans from improving pre-training strategies to adaptation and fine-tuning methods. For instance, a multi-task optimization algorithm has been developed to automatically identify the optimal data fusion strategies, enhancing models’ robustness against cross-modal distribution shifts (Ma et al., 2022). Andonian et al.(Andonian et al., 2022) presents a novel contrastive learning method for pretraining, combining knowledge distillation and soft alignments to improve robustness. Other studies include devising robust adaptation methods. For instance, (Wortsman et al., 2022) shows that fusing weights from zero-shot and fine-tuned CLIPs enhances distributional robustness, with further improvements observed through weight averaging from CLIPs fine-tuned under different hyperparameters.
2.2. Adversarial Robustness
@@
MMLS robustness to adversarial attacks faces unique challenges. Malicious users might use the information from multiple modalities to breach security. Cross-modal adversarial attacks, such as subtle alterations in an image affecting text predictions, further complicate the issue (Zhang et al., 2022).
2.2.1. Multimodal adversarial attack.
@@
Multimodal adversarial attack aims to trick a target MMLS by subtly altering its multimodal inputs. The objective function for multimodal adversarial attacks can generally be formulated as (Tian and Xu, 2021):
(1) $\mathop{\arg\max}_{\{x_{i}^{adv}\}_{i=1}^{n}}\mathcal{L}(\{x_{i}^{adv}\}_{i=1}^{n},y;\theta)\quad s.t.\,||x_{i}^{adv}-x_{i}||_{p}\leq\epsilon_{i},$
@@
where $\{x_{i}\}_{i=1}^{n}$ denotes the input composed of $n$ modalities, $\{x_{i}^{adv}\}_{i=1}^{n}$ are the generated adversarial samples, and $\epsilon_{i}$ is the perturbation budget for the $i$-th modality. The attacker optimizes $\mathcal{L}$ parameterized by $\theta$ to induce model mispredictions within permissible budgets. An effective multimodal adversarial attack needs to properly address the modality gap between multiple input data modalities. (Zhang et al., 2022) studies multimodal attacks on VLMs, discovering that launching unimodal attacks independently often fails. They then propose Co-Attack, which coordinates image and text attacks through a sequential strategy Another challenge lies in executing effective, downstream-agnostic attacks, especially within the prevalent ‘pretrain-then-finetune’ framework for MMLS, as designing specific attacks for a pre-trained encoder targeting each task is impractical. Pioneering work by (Zhou et al., 2023a) proposes a method for generating broadly applicable adversarial noise, ensuring adversarial samples’ significant deviation from decision boundary in the feature space via a topology-deviation-based generative adversarial network.
2.2.2. Enhance adversarial robustness of multimodal learning models.
@@
Current works on improving the adversarial robustness of MMLS center around robust fusion strategy and adversarial training.
@@
Robust Fusion and Alignment. The unique modality alignment process of MMLS, compared to unimodal models, has catalyzed extensive research into enhancing their adversarial robustness. For instance, (Yang et al., 2022) improves MMLS single-source adversarial robustness by designing a fusion strategy that detects inconsistencies across modalities and executes feature fusion to counteract the perturbed modality. A constrastive learning framework has been proposed that captures both intra- and inter-modality dynamics, filtering out noise present in various modalities and thereby yielding more robust representations (Lin and Hu, 2022).
@@
Adversarial Training. Adversarial training enhances robustness effectively, yet its application in MMLS presents challenges, including higher computational demands and scalability issues. To approach these, (Gan et al., 2020) adopts the free adversarial training strategy that obtains parameters’ gradients with almost no extra computational expense. Another challenge lies in improving the generalization of adversarial training procedures within the emerging pretrain-then-finetune paradigm for MMLS. An adversarial pre-training strategy for VLMs is introduced in (Gan et al., 2020), focusing on adversarial training in the embedding space of each modality rather than directly perturbing image pixels or text tokens, effectively enhancing model performance in a task-agnostic manner.
2.3. Benchmark
@@
The assessment of model distributional robustness commonly employs metrics such as Effective Robustness, along with Relative Robustness and the MultiModal Impact Score to measure average performance drops during distributional shifts (Fang et al., 2022; Qiu et al., 2022). Systematic benchmarks are constantly being proposed as well. For instance, (Li et al., 2020) integrates nine VQA datasets to study four generic robustness in VLMs. Recently, benchmarks for large MMLS have emerged, such as the one introduced in (Han et al., 2023). This benchmark assesses GPT-4V’s (Achiam et al., 2023) zero-shot generalization on 13 varied datasets, offering insights into its adaptability to distribution shifts.
2.4. Discussion